OpenStudy (anonymous):

I want to know If my answer is correct. Solve the formula : 4x - 7y = 11. I got.. 4x/7 + 11/7?

OpenStudy (anonymous):

4x-7y=11 7y=4x-11 y=(4/7)x-11/7

OpenStudy (anonymous):

Can that also be written as (4x-11) / 7 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

your welcome